OpenStudy (anonymous):

Need help A ball is thrown directly up in the air from a height of 5 feet with an initial velocity of 60 ft/s. Ignoring air resistance, how long until the ball hits the ground? Use the formula h=-16t^2+60t+5, where h is the height of the ball in feet and t is the time in seconds since it is thrown. Round your answer to the nearest tenth A. 2.1 seconds B. 3.8 seconds C. 2.4 seconds D. 1.8 seconds

OpenStudy (anonymous):

I say it's B not sure tho

OpenStudy (anonymous):

Yep b

OpenStudy (anonymous):

'm gonna give you a short way to do this Let us name our vectors! Anything going up is positive and anything coming down / acting down is negative. u = +60 ft/s g = -9.8 m/s^2 (since gravity acts downward) In this case though, it is m/s^2 and not ft/s^2....Change metres to feet and we will get -32.1522 ft/s^2 s = -5 ft (s is displacement which is vector and that is the direction the ball will be falling when coming down so we have taken into consideration direction). t = what we are looking for We can therefore use the equation: s = ut + 0.5at^2 We can modify ti to get: s = ut + 0.5gt^2 Plug in values: -5 = (60)t + 0.5(-32.1522)t^2 -5 = 60t - 16.0761t^2 FORM INTO A QUADRATIC EXPRESSION: -16.0761t^2 + 60t + 5 = 0 Solve for t by using the quadratic equation: You get t = -0.08 s (1 d.p) and t = 3.8 s TIME CAN NEVER BE NEGATIVE, so we will use the POSITIVE value for the time Hence, the time for trajectory (movement of the ball) is 3.8 seconds